# Turbulent gas-rich discs at high redshift: origin of thick stellar discs through 3D ‘baryon sloshing’

### v1

@@ 4.3 Brownian motion
We separate the phenomenon and consequences of baryon sloshing in the total gravitational potential into two parts, which can both be explored with the Fokker-Planck equation using a probability density function $P(\bm{R}-\bm{R}_{0},t)$, for which $\bm{R}$ is the radius vector from an initial position $\bm{R}_{0}$. First, the baryon potential minimum moves around stochastically with what is effectively damped Brownian motion. We demonstrate this property with a simple model. Secondly, the internal response of the baryons is subject to energy loss and energy gain in different measures due to the changing underlying potential. We treat the internal disc evolution (heating and cooling) separately as a diffusion problem in action space (see Sec. 5.3).
@@
To simulate the track of a particle, Langevin’s equation is used to describe the motion subject to (over-damped) Brownian motion, i.e. $\gamma\bm{\dot{}}{\bm{R}}\approx\bm{F}(\bm{R})+\bm{\eta}(t)$. In our N-body simulations, the damping is mostly due to dynamical friction, i.e. energy loss to the live dark matter halo. To add an extra layer of realism, we computed the baryon motion in a moderately flattened, spherical gravitational potential (${\bm{F}}=-\nabla\Phi_{0}$) that mimics the potential in Fig. 4. This weakly biases each step $d{\bm{R}}$ along the radius vector towards the centre, and compresses the motion in $z$. An example of such a track is shown in Fig. 10 from two perspectives; most realizations capture the essence of Figs. 7 and 8. In the absence of external forces, we can mimic isotropic diffusion in three dimensions using a 3D Gaussian random variate such that ${\bm{\eta}}(t)={\cal G}[\bm{R}(t)-\bm{R}_{0},{\bm{\sigma}}]$ where $|{\bm{\sigma}}|\propto\sqrt{D_{2}t}$. For an ensemble, we find: (i) the first moment is $\langle\eta_{i}(t)\rangle=0$; (ii) the second moment is $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=2{\cal D}_{2}\delta_{ij}\delta(t-t^{\prime})$ where $\delta_{ij}$ is the Kronecker delta and $\delta$ is the Dirac delta function. The latter ensures the covariance of the (Wiener) noise process is uncorrelated. ${\cal D}_{2}$ is a diffusion coefficient that we chose arbitrarily, along with the timesteps, to match the amplitude of motion in kiloparsecs seen in the N-body simulations.
@@
Figure 11: The rms $z$-scale height for pre-existing (old) stars as a function of radius $R$ (in kpc) for each value of $f_{\rm gas}$. The left panel is for no-halo accretion cases; the right panel is for halo accretion models. In all plots, the colour coding indicates the time evolution, as indicated. The scale height of the $\alpha$-rich thick disc today at $R_{0}=8.2$ kpc is indicated by the cross (Bland-Hawthorn & Gerhard, 2016). Note that the vertical scale has a fourfold stretch compared to the horizontal axis. Gas-rich models are more able to reproduce the $\alpha$-rich disc within $1.5-2$ Gyr.
@@
Figure 12: (Top) The rms radial scale length for pre-existing stars as a function of time (in Myr) for each value of $f_{\rm gas}$ indicated in the key. The left panel is for no-halo accretion models; the right panel is for halo accretion models. (Right) The same quantity but for newly formed stars as a function of time (in Myr) for each value of $f_{\rm gas}$ indicated in the key. In all panels, the grey band indicates the radial scale length of the thick disc today, i.e. $R^{T}_{\rm disc}=2.0\pm 0.2$ kpc (Bland-Hawthorn & Gerhard, 2016).
Figure 13: This plot is repeated from Fig. 11 but highlights the different heating phases due to rapid reversals along the sloshing track. The $z-$height is starting to saturate after 1.5 Gyr. Note that the vertical scale has a fourfold stretch compared to the horizontal axis.
@@ 5 Discussion: The consequences of sloshing
Figure 14: Predicted trends at 2 Gyr for the halo accreting $f_{\rm gas}=100\%$ model, representative of the high $f_{\rm gas}$ models in general. (Left) Age-metallicity trend in the simulated $\alpha$-rich disc. (Top right) Age distribution in the simulated $R-z$ plane showing a weak trend in radius but a strong vertical gradient, such that the youngest stars lie near the plane. (Bottom right) Metallicity distribution in the simulated $R-z$ plane showing a weak trend in radius but strong vertical stratification, such that the most metal-rich stars lie near the plane. For the RHS figures, note that the radial trends are weak to non-existent.
@@ 5.1 Disc thickening
The formation of early thick stellar discs has been a challenge for theory and simulations alike. One class of models invoke stars forming in a turbulent, kinematically hot ISM to generate a thick stellar disc, which can be further heated by mergers or interactions with satellites (Stinson et al., 2013; Ma et al., 2017; Bird et al., 2021; van Donkelaar et al., 2022b). However, this effect is not seen in our higher-resolution Nexus simulations. In App. B, we present evidence against disc thickening arising from in situ star formation in a turbulent gas disc. Although star formation and turbulence are still involved, our mechanism is very different.
@@
In Fig. 11, the rate of disc heating increases dramatically with $f_{\rm gas}$. This is shown for pre-existing stars as these are a good tracer of the internal heating. (We present the matching plot for newly formed stars in App. B.) The internal heating continues until the gas fraction drops and the star formation rate follows suit, at which point the internal heating becomes saturated. The high $f_{\rm gas}$ models achieve the $\alpha$-rich thick disc values today. The $f_{\rm gas}=80\%$ and 100% models overshoot the measured values today, but this is likely to have taken place in the early Milky Way. Once the younger, massive thin disc forms, the pre-existing $\alpha$-rich thick disc must have been compressed by the increased density of the $\alpha$-poor disc.
@@
Another consequence of disc heating is the rise of the vertical stellar dispersion $\sigma_{z}$ in both pre-existing and newly formed stars with cosmic time. The relevant figures are presented in our earlier paper (see Figs. 6 and 20 in Bland-Hawthorn et al., 2024). Evidence for quasi-isotropic heating is the roughly preserved ratios $\sigma_{z}/\sigma_{R}$ and $\sigma_{\phi}/\sigma_{R}$ across all radii for the pre-existing stars throughout the simulation. The situation is different for the newly formed stars because their radial distribution exhibits inside-outside growth (Fig. 12), as observed in our movies.
@@
In Fig. 12, the evolution of the radial scale length is presented over the full simulation for the pre-existing (top) and newly formed stars (bottom). To combat noisy data, we chose the rms value as a proxy for the exponential disc scale length $R_{\rm disc}$. For the pre-existing stars, we see the disc grow radially by no more than about 20% of its starting dimension. For newly formed stars, the growth is more dramatic. All halo-accretion models achieve the radial scale length observed for the thick disc today (Bland-Hawthorn & Gerhard, 2016), indicated by the grey band. For the non-accreting models, a very high starting $f_{\rm gas}$ value is required. In contrast, the vertical disc scale height grows by a factor of $1.5-8$ depending on the model (Fig. 11 and App. B ). This statement is not inconsistent with the claim of isotropic heating: the young disc is initially very cold in the vertical direction ($\sigma_{z}\sim 5$ km s-1) compared to the radial direction ($\sigma_{R}\sim 60$ km s-1).
@@ 5.2 Chemical signatures
We provide preliminary results on the predicted stellar chemistry for the early thickened disc (see Tepper-Garcia et al., 2024; Agertz et al., 2021, for details). Some of our “results” are not particularly compelling because, like all simulators, we set up the initial conditions for success. Specifically, the initial star formation rates were very high and this guaranteed the production of high-mass stars and supernovae under starburst conditions. These, in turn, produced $\alpha$-enhanced yields ensuring the first generations of stars are $\alpha$-enriched. Simulators often present their “predicted” old $\alpha$-rich stars as a success of the model, but the first generations tend to emerge as a direct consequence of the initial set-up conditions. Arguably, it is the subsequent generations that are more useful.
In particular, the left panel in Fig. 14 shows a clear trend in [Fe/H] as time passes. This is significant because the thick disc has been described as a snap-frozen, mono-age stellar population (see Sec. 1 and Freeman & Bland-Hawthorn 2002), which is inconsistent with our model. Well defined age-metallicity trends have been discussed extensively in the context of the $\alpha$-rich disc (Haywood et al., 2013b; Hayden et al., 2014; Xiang & Rix, 2022b; Mackereth et al., 2019; Sharma et al., 2021). Most recently, this is precisely what is seen for the $\alpha$-rich population in the GALAH/Gaia (Sahlholdt et al., 2022) and APOGEE/Gaia surveys (Cerqui et al 2025), both in terms of timescale and metallicity range.
This age-metallicity trend is imprinted spatially across the disc. The right panel in Fig. 14 show $R-z$ cross-sections of the thickened stellar disc at 2 Gyr. The upper panel exhibits a clear vertical gradient in stellar age, with the oldest stars furthest from the plane. The lower panel shows a clear vertical stratification in stellar metallicity with the most metal-poor stars furthest from the plane. These general trends (albeit described by different functions of $R$ and $z$) are well established for both the $\alpha$-poor and $\alpha$-rich discs (Bovy et al., 2012a; Mackereth et al., 2017b).
@@ 5.3 Internal heating of the stellar disc
In contemporary discussions of Galactic heating, we tend to focus on either vertical or in-plane processes, or heat energy lost to the dark matter halo. This is because the discussion centres around an observed phenomenon and a specific process, e.g. radial abundance gradients and in-plane migration. But here we stress that baryon sloshing must lead to isotropic heating, at least to first order, such that all Galactic coordinates are affected simultaneously. Figs. 11 and 12 are evidence enough that vertical and in-plane heating occur concurrently in the simulations. More discussion is warranted to emphasize this important point.
Galaxies live in action space -- this is the appropriate framework for discussing internal processes in complex simulations (Binney, 2025). Indeed, the Nexus simulations are initially set up with action-based modelling defined by distribution functions $f_{i}(\bm{J})$ that are analytic functions of the action integrals ${\bm{J}}=(J_{R},J_{\phi},J_{z})$ (Sec. 2). We now provide an argument for why baryon sloshing leads to quasi-isotropic heating of all actions, although the effects are most evident in $J_{z}$ because this property of galaxy discs is more compressed initially. In our simulations, all stars are moving in a time-dependent gravitational potential. Averaged over a dynamical time ($\tau_{\rm dyn}\approx 2\pi/\Omega_{\phi}$), there is an underlying smooth potential that evolves slowly, $\Phi_{0}$. On top of this, there is a fluctuating potential that evolves rapidly, $\Phi_{1}$. The respective distribution functions sum to give $f=f_{0}+f_{1}$. Some of the fluctuations are due to periodic forcing, as we show, but others arise from a jolting, jostling or sloshing action; these variations can be of order the strength of the underlying potential.
For $\Phi_{0}$, the actions are conserved quantities of integrable systems averaged over a dynamical time. Thus a star’s orbital energy can be written as $E={\mathbf{\Omega}}.{\mathbf{J}}$. The radial and vertical actions describe the amount of radial and vertical activity, and the azimuthal action is the angular momentum about the $z$-axis ($J_{\phi}\equiv L_{z}$). In disc galaxies, most stars are born in a roughly 1D action space [^6] In the Nexus simulations, we find this statement to be true. But in cosmological (e.g. FIRE) simulations, which are typically at much lower resolution, some authors find that stars born in the early bursty, gas-rich phase mostly have a thicker vertical distribution and thus significant vertical and radial actions (e.g. Yu et al., 2021; McCluskey et al., 2024). Energy-conserving stars initially on planar, circular orbits tend to become more radial when lifted out of the plane., i.e. in cold, near-circular, near-planar orbits (i.e. large $J_{\phi}$, small $J_{R}$, $J_{z}$). A random perturbation (e.g. jolting action) of the system causes a star to migrate in 3D action space away from its initial conditions.
The Fokker-Planck equation can also be used to describe the time evolution of $f(\mathbf{J})$; it is widely used to model systems influenced by a combination of deterministic and random forces. It traces the evolution of phase space through a running series of equilibria. Let $P(\bm{J},\bm{\Delta})d^{3}\bm{\Delta}$ be the probability per unit time a star with actions $\bm{J}$ migrates to actions in a volume $d^{3}\bm{\Delta}$ centred on $\bm{J}$ (i.e. $\bm{J}\pm\bm{\Delta}/2$).
By analogy with Sec. 4.3, the first two terms of the diffusion equation reduce to (Binney, 2025)
$\frac{\partial f}{\partial t}=-\frac{\partial(f\overline{\Delta}_{i})}{\partial J_{i}}+\frac{1}{2}\frac{\partial^{2}(f\overline{{\Delta_{ij}}^{2}})}{\partial J_{i}\partial J_{j}}$ (25)
where
$\overline{\Delta}_{i}(\bm{J})$ $=$ $\int\Delta_{i}P(\bm{J},\bm{\Delta})\;d^{3}\bm{\Delta}$ (26)
$\overline{{\Delta_{ij}}^{2}}(\bm{J})$ $=$ $\int\Delta_{i}\Delta_{j}P(\bm{J},\bm{\Delta})\;d^{3}\bm{\Delta}.$ (27)
The diffusive flux ${\cal F}$ is therefore
${\cal F}=\bm{D}^{(1)}f+\bm{D}^{(2)}.\frac{\partial f}{\partial\bm{J}}$ (28)
where ${\bm{D}^{(1)}}=\overline{\bm{\Delta}}-\frac{1}{2}\frac{\partial\overline{\bm{\Delta}^{2}}}{\partial\bm{J}}$ and ${\bm{D}^{(2)}}=-\frac{1}{2}\overline{\bm{\Delta}^{2}}$ are the first (vector) and second order (3-tensor) diffusion coefficients. ${\bm{D}^{(1)}}$ is the drift term and captures the energy loss due to dynamical friction, dissipation and external forces. ${\bm{D}^{(2)}}$ is the fluctuating term and captures the energy gain due to noisy processes; it also handles any anisotropic diffusion. Overall, the system’s energy $E={\bm{\Omega}}.{\cal F}$ therefore evolves as
$\frac{dE}{dt}=(2\pi)^{3}\int{\bm{\Omega}}.{\cal F}\;d^{3}\bm{J}$ (29)
Seen in this light, the migration in action space is essentially a random walk at each step due to the fluctuating potential. This may not guarantee an overall energy gain if energy is lost due to, say, heating by dynamical friction of the dark matter halo. To establish whether Eq. 29 leads to a gain or a loss, Binney (2025) describes how ${\bm{D}^{(1)}}$ and ${\bm{D}^{(2)}}$ can be measured directly from the autocorrelation coefficients of the fluctuating DF $f_{1}$ or equivalently the fluctuating potential $\Phi_{1}$, as these are positively correlated. This can be shown to be true using Hamilton’s equation, even in the presence of a time lag, without the need to assume instantaneity from Poisson’s equation. Evidently, the internal heating in action space clearly happens, regardless of how the baryons are being sloshed by Brownian motion. A detailed analysis of this kind is beyond the scope of the present study, but will be addressed in future work.

### v2

@@ 4.3 Brownian motion
We separate the phenomenon and consequences of baryon sloshing in the total gravitational potential into two parts, which can both be explored with the Fokker-Planck equation using a probability density function $P(\bm{R}-\bm{R}_{0},t)$, for which $\bm{R}$ is the radius vector from an initial position $\bm{R}_{0}$. First, the baryon potential minimum moves around stochastically with what is effectively damped Brownian motion. We demonstrate this property with a simple model. Secondly, the internal response of the baryons is subject to energy loss and energy gain in different measures due to the changing underlying potential. In a later paper, we treat the internal disc evolution (heating and cooling) separately as a diffusion problem in action space.
@@
To simulate the track of a particle, Langevin’s equation is used to describe the motion subject to (over-damped) Brownian motion, i.e. $\gamma\bm{\dot{}}{\bm{R}}\approx\bm{F}(\bm{R})+\bm{\eta}(t)$. Here, $\bm{F}$ is the continuously applied force (e.g. galactic potential gradient), $\eta$ supplies a random force term and $\gamma$ is a damping constant. In our N-body simulations, the damping is mostly due to dynamical friction, i.e. energy loss to the live dark matter halo. To add an extra layer of realism, we computed the baryon motion in a moderately flattened, spherical gravitational potential (${\bm{F}}=-\nabla\Phi_{0}$) that mimics the potential in Fig. 4. This weakly biases each step $d{\bm{R}}$ along the radius vector towards the centre, and compresses the motion in $z$. An example of such a track is shown in Fig. 10 from two perspectives; most realizations capture the essence of Figs. 7 and 8. In the absence of external forces, we can mimic isotropic diffusion in three dimensions using a 3D Gaussian random variate such that ${\bm{\eta}}(t)={\cal G}[\bm{R}(t)-\bm{R}_{0},{\bm{\sigma}}]$ where $|{\bm{\sigma}}|\propto\sqrt{D_{2}t}$. For an ensemble, we find: (i) the first moment is $\langle\eta_{i}(t)\rangle=0$; (ii) the second moment is $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=2{\cal D}_{2}\delta_{ij}\delta(t-t^{\prime})$ where $\delta_{ij}$ is the Kronecker delta and $\delta$ is the Dirac delta function. The latter ensures the covariance of the (Wiener) noise process is uncorrelated. ${\cal D}_{2}$ is a diffusion coefficient that we chose arbitrarily, along with the timesteps, to match the amplitude of motion in kiloparsecs seen in the N-body simulations.
@@
Figure 11: The rms $z$-scale height for pre-existing (old) stars as a function of radius $R$ (in kpc) for each value of $f_{\rm gas}$. The left panel is for no-halo accretion cases; the right panel is for halo accretion models. In all plots, the colour coding indicates the time evolution, as indicated. The scale height of the $\alpha$-rich thick disc today at $R_{0}=8.2$ kpc is indicated by the cross (Bland-Hawthorn & Gerhard, 2016). Note that the vertical scale has a fourfold stretch compared to the horizontal axis. Gas-rich models are more able to reproduce the $\alpha$-rich disc within 1.521.5-2 Gyr.
@@
Figure 12: (Top) The radial scale length for pre-existing stars as a function of time (in Myr) for each value of $f_{\rm gas}$ indicated in the key. The left panel is for no-halo accretion models; the right panel is for halo accretion models. (Right) The same quantity but for newly formed stars as a function of time (in Myr) for each value of $f_{\rm gas}$ indicated in the key. In all panels, the grey band indicates the radial scale length of the thick disc today, i.e. $R^{T}_{\rm disc}=2.0\pm 0.2$ kpc (Bland-Hawthorn & Gerhard, 2016).
Figure 13: (Left) The $f_{\rm gas}=100\%$ accreting halo case is repeated from Fig. 11 highlights the different heating phases due to rapid reversals along the sloshing track. In practice, we allow for 1% of the disc mass to be an existing stellar population so that we can trace the impact of the heating.
The $z$-height is starting to saturate after 1.5 Gyr. (Right) This is a repeat of the LHS plot but for stars forming in the gas rather than pre-existing stars. The heating rate is essentially the same. The bottom 10 tracks (100 Myr apart) are removed because of large intrinsic scatter; this is a consequence of small number statistics and early populations being born out of dynamical equilibrium. Note for both plots that the vertical scale has a fourfold stretch compared to the horizontal axis. The cross indicates the vertical scale height of the Milky Way’s $\alpha$-rich disk at the Solar Circle today.
@@ 5 Discussion: The consequences of sloshing
Figure 14: Predicted trends at 2 Gyr ($z\approx 1.5$) after disc onset for the halo accreting $f_{\rm gas}=100\%$ model, representative of the high $f_{\rm gas}$ models in general. (Left) Age-metallicity trend in the simulated $\alpha$-rich disc in the simulated RzR-z plane, i.e. within a vertical slice that is 100 pc thick along the line of sight. (Top right) Age distribution in the same plane showing a weak trend in radius but a stronger vertical gradient, such that the youngest stars lie near the plane. (Bottom right) Metallicity distribution in the same plane showing a weak trend in radius but a stronger vertical stratification, such that the most metal-rich stars lie near the plane. For the RHS figures, representative stellar density contours within the vertical slice are overlaid.
@@ 5.1 Disc thickening
The formation of early thick stellar discs has been a challenge for theory and simulations alike. One class of models invoke stars forming in a turbulent, kinematically hot ISM to generate a thick stellar disc, which can be further heated by mergers or interactions with satellites (Stinson et al., 2013; Ma et al., 2017; Bird et al., 2021; van Donkelaar et al., 2022). However, this effect is not seen in our higher-resolution Nexus simulations. In Sec. 5.4, we present evidence against disc thickening arising from in situ star formation in a turbulent gas disc. Although star formation and turbulence are still involved, our mechanism is very different.
@@
In Fig. 11, the rate of disc heating increases dramatically with $f_{\rm gas}$. This is shown for pre-existing stars as these are a good tracer of the internal heating. (We present the matching plot for newly formed stars in Sec. 5.4.) The internal heating continues until the gas fraction drops and the star formation rate follows suit, at which point the internal heating becomes saturated. The high $f_{\rm gas}$ models achieve the $\alpha$-rich thick disc values today. The $f_{\rm gas}=80\%$ and 100% models overshoot the measured values today, but this is likely to have taken place in the early Milky Way. Once the younger, massive thin disc forms, the pre-existing $\alpha$-rich thick disc must have been compressed by the increased density of the $\alpha$-poor disc.
@@
Another consequence of disc heating is the rise of the vertical stellar dispersion $\sigma_{z}$ in both pre-existing and newly formed stars with cosmic time. The relevant figures are presented in our earlier paper (see Figs. 6 and 20 in Bland-Hawthorn et al., 2024). In Sec. 5.3, we present a simple model for how this heating arises from the fluctuating gravitational potential. Evidence for quasi-isotropic heating is the roughly preserved ratios $\sigma_{z}/\sigma_{R}$ and $\sigma_{\phi}/\sigma_{R}$ across all radii for the pre-existing stars throughout the simulation. Initially, the situation is different for the newly formed stars because their radial distribution exhibits inside-outside growth (Fig. 12), as observed in our movies. But the heating rate of the younger stars matches that of the pre-existing stars once the formed stars extend across the pre-existing disc (see Sec. 5.3).
@@
In Fig. 12, the evolution of the radial scale length is presented over the full simulation for the pre-existing (top) and newly formed stars (bottom). To combat noisy data, we determine the disc scale length $R_{\rm disc}$ from the radius enclosing half the disc’s mass; we divide the latter by 1.67, which is appropriate for a 3D exponential disc. For the pre-existing stars, we see the disc grow radially by no more than about 20% of its starting dimension. For newly formed stars, the growth is more dramatic. All halo-accretion models achieve the radial scale length observed for the thick disc today (Mackereth et al., 2017; Bland-Hawthorn & Gerhard, 2016), indicated by the grey band. For the non-accreting models, a very high starting $f_{\rm gas}$ value is required. In contrast, the vertical disc scale height grows by a factor of 1.581.5-8 depending on the model (Fig. 11 and App. 5.4 ). This statement is not inconsistent with the claim of isotropic heating: the young disc is initially very cold in the vertical direction ($\sigma_{z}\sim 5$ km s-1) compared to the radial direction ($\sigma_{R}\sim 60$ km s-1).
@@ 5.2 Chemical signatures
We provide preliminary results on the predicted stellar chemistry for the early thickened disc (see Tepper-García et al., 2024; Agertz et al., 2021, for details). Some of our “results” are not particularly compelling because, like all simulators, we set up the initial conditions for success. Specifically, the initial star formation rates were very high and this guaranteed the production of high-mass stars and supernovae under starburst conditions. These, in turn, produced $\alpha$-enhanced yields ensuring the first generations of stars are $\alpha$-enriched. This signature remains as some of the best supporting evidence that early discs were gas-rich (Freeman & Bland-Hawthorn, 2002; Bland-Hawthorn & Sharma, 2016). Simulators often present their “predicted” old $\alpha$-rich stars as a success of the model, but the first generations tend to emerge as a direct consequence of the initial set-up conditions. Arguably, it is the subsequent generations that are more useful.
In particular, the left panel in Fig. 14 shows a clear trend in [Fe/H] as time passes. This is significant because the thick disc has been described as a snap-frozen, mono-age stellar population (see Sec. 1 and Freeman & Bland-Hawthorn 2002), which is inconsistent with our model. Well defined age-metallicity trends have been discussed extensively in the context of the $\alpha$-rich disc (Haywood et al., 2013; Hayden et al., 2014; Xiang & Rix, 2022; Mackereth et al., 2019; Sharma et al., 2021). Most recently, this is precisely what is seen for the $\alpha$-rich population in the GALAH/Gaia (Sahlholdt et al., 2022) and APOGEE/Gaia surveys (Cerqui et al., 2025), both in terms of timescale and metallicity range.
The right panel in Fig. 14 show RzR-z cross-sections of the thickened stellar disc at 2 Gyr ($z\approx 1.5$) after the disc formed at $z\approx 3.3$. The upper panel exhibits a clear vertical gradient in stellar age, with the oldest stars furthest from the plane. The lower panel shows a clear vertical stratification in stellar metallicity with the most metal-poor stars furthest from the plane. These general trends (albeit described by different functions of $R$ and $z$), including the flaring at large radius due to the declining disc potential, are well established for both the $\alpha$-poor and $\alpha$-rich discs (Bovy et al., 2012a; Mackereth et al., 2017).
@@ 5.3 Internal heating of the stellar disc
Contemporary discussions of Galactic heating tend to focus on either vertical or in-plane processes, or heat energy lost to the dark matter halo, although 3D disc heating has been discussed (Ida et al., 1993). This is because the discussion centres on an observed phenomenon and a specific process, e.g. radial abundance gradients and in-plane migration. But here we stress that baryon sloshing must lead to isotropic heating, at least to first order, such that all Galactic coordinates are affected simultaneously. Figs. 11 and 12 are evidence enough that vertical and in-plane heating occur concurrently in the simulations. In a later paper, we return to a detailed discussion of 3D heating using a Fokker-Planck diffusion analysis in action space.
Instead, we use a simple model for how vertical heating of the stellar disc arises from sloshing in a time-dependent background potential. The connection between fluctuating gravitational fields and internal dissipation or heating is a classical problem in statistical physics (Green, 1954; Kubo et al., 1957), and has also been examined in astrophysical dynamics (e.g. Nelson & Tremaine, 1999; Weinberg, 2001). To build intuition, consider a galactic disc oscillating vertically within an external potential. If the fluctuation $\delta\Phi$ has a zero mean around the equilibrium position ($z=0$), the equation of motion for a stellar mass $m$ is $m\dot{v}_{z}=F(z,t)=-\nabla\delta\Phi$. For simplicity, we set $m=1$ since the mass cancels out in the dimensionless form of the analysis. For a stochastic process, the diffusion coefficient is defined as
${\cal D}$ $=$ $\frac{1}{2}\frac{d}{dt}\langle(v_{z}(t)-v_{z}(0))^{2}\rangle$ (25)
$=$ $\frac{d}{dt}\int_{0}^{t}(t-\tau)\>{\cal C}(\tau)\>d\tau$ (26)
where the force autocorrelation function is ${\cal C}(\tau)=\langle F(0)F(\tau)\rangle$. As $t$ goes to infinity, the diffusion coefficient simplifies to
${\cal D}=\int_{0}^{\infty}{\cal C}(\tau)\>d\tau\approx\langle F^{2}\rangle\tau_{\rm corr}$ (27)
assuming $\cal C$} decays over a correlation time $\tau_{\rm corr}$ and has a variance $\langle F^{2}\rangle$. The baryon sloshing decays over timescales of 400-700 Myr for the different models as the gas is converted into stars. We approximate $\langle F^{2}\rangle=(\delta\Phi_{z}/Z)^{2}$ where ZZ is the vertical distance over which the potential is varying. For $\Phi_{z}\approx\sigma_{z}^{2}$ and $\tau_{\rm dyn}\approx Z/\sigma_{z}$, assuming the simulation time $T_{o}$ (2 Gyr) is much longer than $\tau_{\rm dyn}$ ($\approx 600/2\pi$ Myr), we can approximate the change in the stellar velocity dispersion to be
$\frac{\delta\sigma_{z}^{2}}{\sigma_{z}^{2}}\approx\frac{2}{3}\left(\frac{\delta\Phi_{z}}{\Phi_{z}}\right)^{2}\frac{\tau_{\rm corr}\>T_{o}}{\tau^{2}_{\rm dyn}}.$ (28)
as a function of the change in the underlying gravitational potential.
The clearest demonstration of how stellar populations heat with time is presented in Bland-Hawthorn et al. (2024, Fig. 6), drawn from the same data used in Fig. 11. For the highest gas-fraction models, the pre-existing stars increase their dispersions by factors of 1.521.5-2, requiring gravitational fluctuations of $10-20\%$. The young stars, at least those born in a dynamically stable disc, have a colder intrinsic dispersion ($\sigma_{z}\approx 6-8$ km s-1): these increase their dispersions by factors of 343-4 requiring fluctuations of 50%. This is an order of magnitude too high to be explained by the underlying dark matter potential or by stars being born out of equilibrium. Instead, we believe the extra heating arises from the large $\delta\Phi_{z}$ fluctuations due to the unbinding of the disc. We refer the reader to Fig. 3 and the discussion in Sec. 3.1 where $30-50\%$ fluctuations are seen in the most extreme cases. How the unbinding works with the sloshing action is unclear and deserves further study. This work has started and will be reported on in a future paper.
Figure 15: (Left) Where new stars are born in time as a function of vertical height $|z|.$ The left-hand side (LHS) panels look very similar for all $f_{\rm gas}$, quite unlike the RHS panels. (Right) The rms $z$-scale height for newborn stars as a function of radius $R$ (in kpc) for each value of $f_{\rm gas}$ for the halo accretion cases. In all plots, the colour coding indicates the time evolution, as indicated. The bottom 10 tracks (100 Myr apart) are removed because of large intrinsic scatter; this is a consequence of small number statistics and early populations being born out of dynamical equilibrium.
The scale height of the $\alpha$-rich thick disc today at $R=8$ kpc are indicated by the blue star (Bland-Hawthorn & Gerhard, 2016). Note that the vertical scale has a fourfold stretch compared to the horizontal axis. The gas-rich models are able to reproduce the thickness of the $\alpha$-rich disc within 1.521.5-2 Gyr.
5.4 The case against thick discs ($|z|\approx 1$ kpc) from in-situ star formation
In Secs. 5.1 and 5.3, we presented evidence for disc heating being a gradual accumulation of energy from baryon sloshing within the dark matter potential. Star formation feedback and turbulence are implicated in the thickening of discs, but not from stars being born in situ far from the cold molecular disc. This idea first arose from non-physical, sticky particle simulations where there is no equation of state, and therefore no pressure balance, or even feedback from star formation processes to support the gas. In Fig. 15, most stars in dynamical equilibrium are born close to the plane within the scale height of the cold gas. The remaining 20-30% of stars are injected under non-equilibrium conditions. The LHS figures are similar in their temporal behaviour for all $f_{\rm gas}$, unlike the RHS figures. In the RHS panel, the disc heating rate for formed stars in time increases rapidly with $f_{\rm gas}$, all starting from a low base set by the maximum dispersion allowed to turbulent-pressure supported molecular gas ($\leq 20-30$ km s-1; Jeffreson et al., 2022). This is because the Brownian motion arising from baryon sloshing heats the stellar disc essentially isotropically, and the amplitude of this motion is a strong function of $f_{\rm gas}$. Eventually, as the younger stellar component settles, its heating rate emerges from the noise and matches the pre-existing population. This is evident by matching the different curves one for one in Fig. 13, starting with the oldest tracks and moving backwards in time.